# Distance-Domain Degrees of Freedom in Near-Field Region

### v3

@@ V-B Approximation Accuracy for Distance-Domain DoF in the Non-Broadside Case
Our analytical DoF expression is derived under two standard assumptions: (i) a phase model based on the Fresnel approximation; and (ii) a uniform-amplitude model that neglects polarization, and elementwise Tx–Rx distance variation. To assess the impact of these assumptions on the distance-domain DoF, we compare three channel models:
  • Fresnel phase & uniform amplitude: the model used to derive the closed-form DoF for the non-broadside case;
  • Exact phase & uniform amplitude: removes the Fresnel phase approximation while keeping amplitude uniform;
  • Exact phase & exact amplitude: further incorporates polarization and elementwise distance variation in amplitude [lu2021communicating].
For a propagation direction $\mathbf{k}_{k}$ and a Tx element at $\mathbf{p}\!\in\!\mathcal{P}$, the Fresnel expansion introduces a residual phase error whose worst-case bound over the Tx aperture is
$e_{\text{phase}}\;\leq\;\frac{2\pi}{\lambda}\,\frac{\bigl(\mathbf{k}_{k}^{\mathsf{T}}\mathbf{p}_{\max}\bigr)\Bigl(\|\mathbf{p}_{\max}\|^{2}-\bigl(\mathbf{k}_{k}^{\mathsf{T}}\mathbf{p}_{\max}\bigr)^{2}\Bigr)}{2\,r_{\min}^{2}},$ (52)
where $\mathbf{p}_{\max}\!\in\!\mathcal{P}$ attains the maximum magnitude of $\mathbf{k}_{k}^{\mathsf{T}}\mathbf{p}$ over the aperture and $r_{\min}$ is the minimum Tx–Rx distance. The Fresnel distance is the range at which this bound falls below a prescribed maximum phase error (e.g., $\pi/4$); beyond this distance, the Fresnel model is accurate for DoF analysis.
We compute the eigenvalue spectra of the three channel models via numerical eigenvalue decomposition. The Tx array is a square aperture of size $160\lambda\times 160\lambda$ (so the Tx radius satisfies $p_{\min},p_{\max}]=[0\lambda,\,80\sqrt{2}\lambda]$). The Rx array is placed at an azimuth $\phi=75^{\circ}$ with fixed $r_{\max}=4000\lambda$, while $r_{\min}$ varies from $300\lambda$ to $2000\lambda$. The polarization is taken along the $z$-axis. For these parameters, the Fresnel distance for a maximum phase error of $\pi/4$ is approximately $1060\lambda$.
@@
Fig. 10 (top) reports the maximal Fresnel phase error versus $r_{\min}$; the bottom panel compares the normalized eigenvalue distributions for the three models alongside the analytical DoF. When $r_{\min}=2000\lambda$, all three spectra coincide; even at $r_{\min}=1000\lambda$ (near the Fresnel distance), deviations are negligible. This indicates that for sufficiently large Tx–Rx separation, both amplitude non-idealities (polarization and distance variation) and Fresnel phase errors have a minimal impact on the DoF. When $r_{\min}\leq 500\lambda$ (i.e., within the Fresnel region and with distance comparable to the Tx array’s size), noticeable differences arise between the three spectra. Nevertheless, the eigenvalue cutoff index—the point where eigenvalues drop sharply—remains consistent across models and stays within $\approx 10\%$ of the analytical DoF prediction. These simulations suggest that, even when the underlying assumptions are partially violated, the analytical DoF remains a robust predictor of the distance-domain DoF in non-broadside near-field conditions.
In conclusion, our simulations verify that when $r_{\min}$ exceeds the Fresnel distance (corresponding to a maximum phase error of $\pi/4$), the analytical DoF perfectly matches the numerical results from the exact models. Furthermore, even when $r_{\min}$ falls below this Fresnel distance, the deviation between the analytical prediction and the exact models remains negligible. This confirms that the derived DoF expression is robust against polarization effects and amplitude variations, even when the strict Fresnel conditions are partially relaxed.
VI Extensions to Distance-Domain Degrees of Freedom with Modular Array
In this section, we extend the analysis to the more general case in which the Tx array $\mathcal{P}$ consists of multiple non-overlapping sub-arrays.
Figure 11: Geometric setup of the Tx array $\mathcal{P}$ with multiple sub-arrays and the Rx broadside linear array $\mathcal{Q}$.
@@ VI-A Analytical Framework for Spatial DoF of Modular Arrays
Consider a Tx array $\mathcal{P}$ consisting of $N$ non-overlapping continuous sub-arrays, denoted as $\mathcal{P}^{(1)},\;\mathcal{P}^{(2)},\;\dots,\;\mathcal{P}^{(N)}$, where each sub-array $\mathcal{P}^{(n)}$ has their own extreme distances given by:
$p^{(n)}_{\min}=\min_{\mathbf{p}\in\mathcal{P}^{(n)}}\|\mathbf{p}\|,\ p^{(n)}_{\max}=\min_{\mathbf{p}\in\mathcal{P}^{(n)}}\|\mathbf{p}\|,$ (53)
Then, each sub-array $\mathcal{P}^{(n)}$ has an associated Fourier transform (cf. Lemma 2) expressed as
$\tilde{g}^{(n)}(\xi)=\text{sinc}(2\xi)*\tilde{g}_{0}^{(n)}(\xi).$ (54)
where $\tilde{g}_{0}^{(n)}(\xi)$ is given by:
$\tilde{g}_{0}^{(n)}(\xi)=\iint_{\mathbf{p}\in\mathcal{P}^{(n)}}\delta\!\Biggl(\xi+\frac{\|\mathbf{p}\|^{2}}{2\lambda}\Bigl(\frac{1}{r_{\min}}-\frac{1}{r_{\max}}\Bigr)\Biggr)d\mathbf{p}.$ (55)
Using Theorem 1, the main lobe of $\tilde{g}_{0}^{(n)}(\xi)$ lies within the interval:
$\Omega^{(n)}=\left[-\frac{\left(p^{(n)}_{\max}\right)^{2}\Bigl(r^{-1}_{\min}-r^{-1}_{\max}\Bigr)}{2\lambda},\;-\frac{\left(p^{(n)}_{\min}\right)^{2}\Bigl(r^{-1}_{\min}-r^{-1}_{\max}\Bigr)}{2\lambda}\right].$ (56)
Because the sub-arrays are disjoint, their contributions add linearly:
$\tilde{g}(\xi)=\tilde{g}_{0}(\xi)*\text{sinc}(2\xi)$
$=\iint_{\mathbf{p}\in\cup_{n=1}^{N}\mathcal{P}^{(n)}}\delta\!\Biggl(\xi+\frac{\|\mathbf{p}\|^{2}}{2\lambda}\Bigl(r^{-1}_{\min}-r^{-1}_{\max}\Bigr)\Biggr)d\mathbf{p}*\text{sinc}(2\xi)$
$=\sum_{n=1}^{N}\iint_{\mathbf{p}\in\mathcal{P}^{(n)}}\delta\!\Biggl(\xi+\frac{\|\mathbf{p}\|^{2}}{2\lambda}\Bigl(r^{-1}_{\min}-r^{-1}_{\max}\Bigr)\Biggr)d\mathbf{p}*\text{sinc}(2\xi)$
$=\sum_{n=1}^{N}\tilde{g}^{(n)}(\xi).$ (57)
Hence, the overall main-lobe support of $\tilde{g}(\xi)$ is the union of the sub-intervals:
$\Omega=\cup_{n=1}^{N}\Omega^{(n)}$ (58)
The effective bandwidth $\mathrm{BW}(\Omega)$ is therefore obtained by summing the lengths of the $\Omega^{(n)}$ intervals and subtracting any overlaps; equivalently, it is determined by how the radial ranges $p^{(n)}_{\min},\,p^{(n)}_{\max}$] overlap across sub-apertures.
@@
Figure 12: Example: Tx array $\mathcal{P}$ with multiple sub-linear arrays and the Rx broadside linear array $\mathcal{Q}$.
Example (three-subarray ULA).
For simplicity of explaining the concept, let us consider the symmetric modular ULA with 3 sub-arrays as shown in Fig. 12. From (56): (i) $\Omega^{(1)}$ and $\Omega^{(3)}$ completely overlap (the two outer sub-apertures share the same radial extremes), while (ii) $\Omega^{(2)}$ (the central sub-aperture) is disjoint from both. Thus,
$\mathrm{BW}(\Omega)\;=\;\mathrm{BW}\big(\Omega^{(1)}\big)\;+\;\mathrm{BW}\big(\Omega^{(2)}\big),$ (59)
and removing $\mathcal{P}^{(3)}$ does not change $\mathrm{BW}(\Omega)$ (hence, the DoF).
To illustrate, consider $r_{\min},r_{\max}]=[100\lambda,\,200\lambda]$, $p_{\min}^{(1)},p_{\max}^{(1)}]=[p_{\min}^{(3)},p_{\max}^{(3)}]=[80\lambda,\,100\lambda]$, and $p_{\min}^{(2)},p_{\max}^{(2)}]=[0,\,40\lambda]$. Figure 13 shows that the eigenvalue distribution for (i) all three sub-arrays and (ii) only the first two sub-arrays are essentially identical and match the analytical DoF of $13$, calculated by using (59).
@@
Figure 13: Eigenvalue distribution for the 3-subarray versus the 2-subarray configuration (final sub-array removed) in the example from Fig. 12.
@@ VI-B1 Distance-domain DoF under physical length constraint
Figure 14: Eigenvalue distributions of LoS channel between $\mathcal{P}$ and $\mathcal{Q}$ where the Tx array $\mathcal{P}$ is either the linear array or the symmetric modular linear array (Same total length in all cases, and the linear array $\mathcal{P}$ with $r_{\min}=200\lambda$ and $r_{\max}=2000\lambda$).
Firstly, let us compare a conventional linear array and a symmetric modular linear array for Tx array, where the total physical array length is maintained. For a symmetric two-module array with extreme ends given by $[a,b]$ and [b,a][-b,-a], the main-lobes of $\Omega^{(1)}$ and $\Omega^{(2)}$ are identical and overlaps each other. Hence, only one sub-array contributes to the distance-domain DoF of the whole array, and the DoF is given by:
$\frac{b^{2}-a^{2}}{2\lambda}\left(r^{-1}_{\min}-r^{-1}_{\max}\right)+O(1).$ (60)
By defining L=baL=b-a as the length of each sub-array, the DoF can also be written as
$\frac{L(a+b)}{2\lambda}\left(r^{-1}_{\min}-r^{-1}_{\max}\right)+O(1).$ (61)
For this special case, one can see that $\frac{(a+b)}{2}\left(r^{-1}_{\min}-r^{-1}_{\max}\right)$ is approximately the angle subtended by the Rx array seen from the center of the linear Tx sub-array $\mathcal{P}^{(1)}$. Hence, the distance-domain DoF in this special case is analogous to the angular-domain DoF derived for the linear array, as discussed in [poon2005degrees].
As shown by (61), shows that for a fixed sub-array length $L$, increasing the extreme edges (i.e., increasing both $a$ and $b$) expands the distance-domain DoF. Intuitively, increasing the extreme edges expands the angle seen from the center of the sub-array, hence increasing the DoF [poon2005degrees]. This is illustrated in Fig. 14, which suggests that a modular array can exploit additional spatial resources in the distance domain compared to a single-piece array with the same overall physical length.
@@ VI-B2 Spatial DoF under extreme edge constraint
Figure 15: Eigenvalue distributions of LoS channel where the Tx array is either the linear array or the symmetric modular linear array (All outer edges have the same size, and the Rx array satisfies $r_{\min}=200\lambda$ and $r_{\max}=2000\lambda$).
Furthermore, consider the scenario where both a Tx linear array and a Tx modular linear array are constrained to the interval [L,L][-L,L]. In the modular array, a hole of total length $2\alpha L$ ($0<\alpha<1$) is introduced, leading to an effective array length of $2(1-\alpha)L$. Under the condition of a fixed gap’s length, the distance-domain DoF is maximized when the gap is centered. For this case, the distance-domain DoF is given by
$\frac{L^{2}(1-\alpha^{2})}{2\lambda}\left(r^{-1}_{\min}-r^{-1}_{\max}\right)+O(1).$ (62)
@@
Thus, although the modular array has a shorter effective length (by a factor of $1-\alpha$) compared to the full continuous array, the loss in spatial DoF is only of order $\alpha^{2}$. For example, as seen in Fig. 15, a modular array with a less than $40\%$ reduction in length results in only about a roughly $16\%$ loss in spatial DoF. This comparison suggests that removing the central area of the Tx array to some extent only slightly reduces the distance-domain DoF. Hence, this highlights the advantage of modular arrays in providing similar spatial multiplexing capabilities with reduced physical area.
VII Conclusion

### v4

@@ V-B Approximation Accuracy for Distance-Domain DoF in the Non-Broadside Case
Our analytical DoF expression relies on two standard simplifications: a Fresnel phase approximation and a uniform-amplitude model (which neglects polarization and element-wise distance variations). To assess the robustness of these assumptions, we compare three channel models: (i) our baseline Fresnel phase & uniform amplitude; (ii) exact phase & uniform amplitude; and (iii) exact phase & exact amplitude, which incorporates full polarization and distance-dependent path loss [lu2021communicating].
For a propagation direction $\mathbf{k}_{k}$ and a Tx element at $\mathbf{p}\!\in\!\mathcal{P}$, the worst-case residual phase error introduced by the Fresnel expansion over the Tx aperture is bounded by:
$e_{\text{phase}}\;\leq\;\frac{2\pi}{\lambda}\,\frac{\bigl(\mathbf{k}_{k}^{\mathsf{T}}\mathbf{p}_{\max}\bigr)\Bigl(\|\mathbf{p}_{\max}\|^{2}-\bigl(\mathbf{k}_{k}^{\mathsf{T}}\mathbf{p}_{\max}\bigr)^{2}\Bigr)}{2\,r_{\min}^{2}}.$ (47)
The Fresnel distance is defined as the minimum range $r_{\min}$ at which this maximum error falls below a prescribed threshold (e.g., $\pi/4$).
We numerically compute the eigenvalue spectra for these three models using a $160\lambda\times 160\lambda$ square Tx array (where $p_{\max}=80\sqrt{2}\lambda$). The $z$-polarized Rx array is positioned at an azimuth $\phi=75^{\circ}$ with a fixed $r_{\max}=4000\lambda$. For this specific configuration, the Fresnel distance for a $\pi/4$ maximum phase error evaluates to approximately $1060\lambda$.
@@
As illustrated in Fig. 10, when the Tx–Rx separation meets or exceeds the Fresnel distance (e.g., $r_{\min}\geq 1000\lambda$), the spectra of all three models perfectly coincide with our analytical prediction. This confirms that amplitude non-idealities and phase errors are utterly negligible in this regime. More importantly, even when $r_{\min}$ falls deep within the Fresnel region (e.g., $r_{\min}\leq 500\lambda$), where minor spectral deviations naturally arise, the eigenvalue cutoff index remains highly consistent across all models—deviating by less than $10\%$ from our analytical prediction. This confirms that our derived closed-form expression is highly robust against amplitude variations, successfully predicting the distance-domain DoF even when strict Fresnel conditions are relaxed.
VI Extensions to Distance-Domain Degrees of Freedom with Modular Arrays
In this section, we extend our fundamental analysis to a more generalized architecture where the Tx array $\mathcal{P}$ consists of multiple non-overlapping sub-arrays.
Figure 11: Geometric setup of the Tx array $\mathcal{P}$ comprising multiple sub-arrays and the broadside Rx linear array $\mathcal{Q}$.
@@ VI-A Analytical Framework for Spatial DoF of Modular Arrays
Consider a Tx array partitioned into $N$ disjoint continuous sub-arrays, $\mathcal{P}=\bigcup_{n=1}^{N}\mathcal{P}^{(n)}$. The extreme radial distances for each individual sub-array $\mathcal{P}^{(n)}$ are defined as:
$p^{(n)}_{\min}=\min_{\mathbf{p}\in\mathcal{P}^{(n)}}\|\mathbf{p}\|,\quad p^{(n)}_{\max}=\max_{\mathbf{p}\in\mathcal{P}^{(n)}}\|\mathbf{p}\|.$ (48)
Because the Fourier transform (cf. Lemma 2) is a linear operator, the total spatial spectrum $\tilde{g}(\xi)$ is simply the superposition of the spectra from each sub-array:
$\tilde{g}(\xi)=\sum_{n=1}^{N}\tilde{g}^{(n)}(\xi),$ (49)
Consequently, based on Theorem 1, the overall main-lobe support $\Omega$ of the modular array is the direct union of the individual sub-intervals:
$\Omega=\bigcup_{n=1}^{N}\Omega^{(n)},$ (50)
where each sub-interval is bounded by its respective radial extremes:
$\Omega^{(n)}=\left[-\frac{(p^{(n)}_{\max})^{2}(r_{\min}^{-1}-r_{\max}^{-1})}{2\lambda},\;-\frac{(p^{(n)}_{\min})^{2}(r_{\min}^{-1}-r_{\max}^{-1})}{2\lambda}\right].$ (51)
The total effective bandwidth $\mathcal{B}(\Omega)$, which directly dictates the distance-domain DoF, is the Lebesgue measure of this combined support. Crucially, $\mathcal{B}(\Omega)$ is determined entirely by how the radial intervals $p^{(n)}_{\min},\,p^{(n)}_{\max}$] overlap across the $N$ sub-apertures.
@@
Figure 12: Example: Symmetric Tx array $\mathcal{P}$ composed of three sub-linear arrays.
Example (Three-Subarray ULA)
To illustrate this spatial redundancy, consider the symmetric modular ULA composed of three sub-arrays depicted in Fig. 12. Because the two outer sub-arrays ($\mathcal{P}^{(1)}$ and $\mathcal{P}^{(3)}$) are symmetric with respect to the origin, they share the exact same radial extremes. Thus, their spectral supports perfectly overlap: $\Omega^{(1)}=\Omega^{(3)}$. Assuming the central sub-array $\mathcal{P}^{(2)}$ occupies a distinctly smaller radial span, it remains disjoint from the outer supports. The total effective bandwidth evaluates to:
$\mathcal{B}(\Omega)\;=\;\mathcal{B}\big(\Omega^{(1)}\big)\;+\;\mathcal{B}\big(\Omega^{(2)}\big).$ (52)
Remarkably, because $\Omega^{(3)}$ is entirely redundant, removing the sub-array $\mathcal{P}^{(3)}$ does not alter the total bandwidth $\mathcal{B}(\Omega)$, meaning it does not reduce the available distance-domain DoF.
To verify this numerically, we simulate a setup with $r_{\min},r_{\max}]=[100\lambda,\,200\lambda]$. The outer sub-arrays span $80\lambda,\,100\lambda]$, and the inner sub-array spans $0,\,40\lambda]$. As shown in Fig. 13, the eigenvalue distributions for the full 3-subarray configuration and the truncated 2-subarray configuration (with $\mathcal{P}^{(3)}$ removed) are essentially identical. Both strictly align with the analytical DoF of $13$ predicted by (52).
@@
Figure 13: Eigenvalue distribution for the 3-subarray versus the 2-subarray configuration (final sub-array removed) from the example in Fig. 12.
@@ VI-B1 Distance-domain DoF under physical length constraint
Figure 14: Element positions of four array configurations in the xz-plane (top panels) and the corresponding eigenvalue distributions (bottom panel) of the LoS channel between $\mathcal{P}$ and $\mathcal{Q}$. The Tx array $\mathcal{P}$ is either a single-piece linear array or a symmetric modular linear array. Both configurations share the same total active physical length $L$, and the Rx array $\mathcal{Q}$ spans from $r_{\min}=200\lambda$ to $r_{\max}=2000\lambda$. The vertical dotted lines represent the analytical DoF in (39).
Firstly, we compare a conventional single-piece linear array and a symmetric modular linear array, maintaining the same total active physical length $L$. For a symmetric two-module array separated by a central gap of size $G$, the sub-arrays span the intervals [(L+G)/2,G/2][-(L+G)/2,-G/2] and $[G/2,(L+G)/2]$. Because the main-lobes of $\Omega^{(1)}$ and $\Omega^{(2)}$ are identical and perfectly overlap, only one sub-array determines the distance-domain DoF of the entire array, yielding:
$\frac{L(L+2G)}{8\lambda}\left(r^{-1}_{\min}-r^{-1}_{\max}\right)+O(1).$ (53)
Equation (53) shows that for a fixed active array length $L$, increasing the gap size $G$ strictly expands the distance-domain DoF.
Figure 15: Geometric comparison of the subtended angles between a single-piece Tx array (left) and a modular Tx array with a central gap (right). Due to structural symmetry, only the right half of each Tx configuration is depicted.
To intuitively explain this gain, we apply the signal space approach [poon2005degrees], where spatial DoF is proportional to the active array length $L$ and the difference in directional cosines of the receiver boundaries: $\text{DoF}\propto L(\cos\Theta_{1}-\cos\Theta_{2})=2L\sin(\frac{\Theta_{2}-\Theta_{1}}{2})\sin(\frac{\Theta_{2}+\Theta_{1}}{2})$. Here, $\Theta_{1}$ and $\Theta_{2}$ denote the extreme angles of the Rx array observed from the Tx elements.
As illustrated in Fig. 15, laterally shifting the active elements outward by introducing a gap $G$ induces two competing geometric effects:
  • Negligible decrease in $\sin(\frac{\Theta_{2}+\Theta_{1}}{2})$: Since the near-field Rx array distances are much greater than the lateral shift, the angles remain close to $\pi/2$. Consequently, this sine term is nearly $1$, and the lateral shift barely reduces it.
  • Significant increase in $\sin(\frac{\Theta_{2}-\Theta_{1}}{2})$: Moving the Tx elements away from the collinear axis dramatically expands the angular spread $(\Theta_{2}-\Theta_{1})$ subtended by the deep longitudinal Rx array.
Since the substantial growth of the difference term far outpaces the marginal reduction in the sum term, the overall distance-domain DoF strictly increases. This theoretical gain is confirmed numerically in Fig. 14, demonstrating that a modular array can exploit additional spatial resources compared to a single-piece array of identical active physical length.
@@ VI-B2 Spatial DoF under extreme edge constraint
Figure 16: Element positions of four array configurations in the xz-plane (top panels) and the corresponding eigenvalue distributions (bottom panel) of the LoS channel where the Tx array is either a single-piece linear array or a symmetric modular linear array. Both configurations are constrained within the same extreme outer edges, and the Rx array spans from $r_{\min}=200\lambda$ to $r_{\max}=2000\lambda$.
Furthermore, consider a scenario where both a single-piece Tx array and a modular Tx array are strictly bounded within the extreme edges [L/2,L/2][-L/2,L/2] (representing a maximum physical footprint of $L$). For the modular array, a central hole of length $\alpha L$ ($0<\alpha<1$) is introduced, reducing the effective active length to $(1-\alpha)L$. Under these strict boundary constraints, a centered gap maximizes the distance-domain DoF, evaluating to:
$\frac{L^{2}(1-\alpha^{2})}{8\lambda}\left(r^{-1}_{\min}-r^{-1}_{\max}\right)+O(1).$ (54)
@@
Thus, although the modular array’s effective active length is reduced by a factor of $1-\alpha$, the resulting loss in spatial DoF is only of order $\alpha^{2}$. For example, as seen in Fig. 16, a modular array with a $40\%$ reduction in active length ($\alpha=0.4$) experiences only a roughly $16\%$ loss in spatial DoF. This confirms that the central region of the Tx array contributes minimally to the distance-domain DoF, highlighting a key advantage of modular arrays: they achieve comparable spatial multiplexing capabilities while significantly reducing the required hardware area.
VII Discussion: Limitations and Future Directions
While the proposed analytical framework provides a closed-form expression and valuable geometric insights into the distance-domain DoF, it is based on several idealized assumptions. These limitations inherently define the scope of our current contribution and establish a clear roadmap for future extensions:
  • Continuous Aperture Idealization vs. Discrete Arrays: This model assumes an infinite number of elements with infinitesimal element spacing to derive an upper bound for the distance-domain DoF. While mathematically convenient and insightful, this idealized framework is not directly implementable. To practically realize the distance-domain DoF, specific sampling strategies for both the transmitters and receivers are required, which will be addressed in our future work in [duong2025sampling].
  • LoS-Dominant Environments vs. Multipath Scattering: The analysis is strictly restricted to the Line-of-Sight (LoS) component. While LoS models are highly valid in sparse-scattering environments such as mmWave and THz frequency bands, they do not capture the full channel complexity of rich scattering environments. Future work can incorporate Non-Line-of-Sight (NLoS) components to evaluate wireless channels with rich scattering, which typically occur at lower frequency carriers.
  • 1D Observation Region vs. 3D observation region: In this current work, we assume a 1D observation region to effectively isolate the angular-domain DoF and mathematically extract the distance-domain DoF. Such geometry corresponds to collinear users, which represents a highly specific multi-user topology. Future research can generalize this framework to 2D or 3D user distributions, enabling the joint optimization of distance- and angular-domain multiplexing.
VIII Conclusion